OpenStudy (anonymous):

Differentiate the Functions. f(u)= (ln(u)/1+ln(2u))

OpenStudy (inkyvoyd):

is this ln(u)/1?

OpenStudy (inkyvoyd):

cause ln(u)/1=ln(u)...

OpenStudy (anonymous):

f(u)=ln(u)/(1+ln(2u))

OpenStudy (inkyvoyd):

okay. for these problems I'd probably brute force it out with the quotient rule, but there maybe a better way...

OpenStudy (debbieg):

I was going to say quotient rule, too. YOu could rewrite the den'r as 1 + ln2 +ln(u) for ease in the differentiation step. Avoids the chain rule.

OpenStudy (inkyvoyd):

lol, forgot the rules for logarithms :S

OpenStudy (debbieg):

Oh right, lol... you had me convinced, but yeah, the product rule, not the exponent thing, applies here. :)

OpenStudy (inkyvoyd):

so glad I'm out of calculus hehe - no loggy log laws in discrete

OpenStudy (anonymous):

so can someone work it out for me because I keep trying and am not getting the right answer

OpenStudy (debbieg):

How about if you show us what you did, and we can tell you where you're going wrong.

OpenStudy (anonymous):

ill try again later. I dont feel like typing it all out